OpenStudy (anonymous):

Find the difference: (-7xᶾ + x² + 8) – (6x² – 7x + 8)

OpenStudy (anonymous):

\[-7x^3 +x^2 +8-6x^2+7x-8=-7x^3-5x^2=-(7x^3+5x^2)\]

OpenStudy (anonymous):

\[=-x^2(7x+5)\]

OpenStudy (anonymous):

thank you